# Improved Off-policy Reinforcement Learning in Biological Sequence Design

### v1

@@ 5 Related work
Designing biological sequences using machine learning methods is widely studied. Bayesian optimization methods (Mockus, 2005; Belanger et al., 2019; Zhang et al., 2022) exploit posterior inference over newly acquired data points to update a Bayesian proxy model that can measure useful uncertainty. However, these methods usually suffer from scalability issues due to the complexity of the Gaussian process (GP) kernel (Belanger et al., 2019) or the difficulty of sampling from an intractable posterior (Zhang et al., 2022).
@@
Offline model-based optimization (MBO) (Kumar & Levine, 2020; Trabucco et al., 2021; Yu et al., 2021; Chen et al., 2022; Kim et al., 2023; Chen et al., 2023a; Yun et al., 2024) also addresses the design of biological sequences using offline datasets only, which can be highly efficient because they do not require oracle queries. These approaches have reported meaningful findings, such as the conservative requirements on proxy models since proxy models tend to give high rewards on unseen samples (Trabucco et al., 2021; Yu et al., 2021; Yuan et al., 2023; Chen et al., 2023b). This supports our approach of adaptive conservatism in the search process. Surana et al. (2024) recently noted that offline design and existing benchmarks are insufficient to reflect biological reliability, indicating that settings without additional Oracle queries might be too idealistic.
@@
Reinforcement learning methods, such as DyNA PPO (Angermueller et al., 2020) and GFlowNets (Bengio et al., 2021; Jain et al., ; Jain et al., 2023; Hernández-García et al., 2024), and sampling with generative models (Brookes & Listgarten, 2018; Brookes et al., 2019; Das et al., 2021; Song & Li, 2023) aim to search the biological sequence space using a sequential decision-making process with a policy, starting from scratch. Similarly, sampling with generative models (Brookes & Listgarten, 2018; Brookes et al., 2019; Song & Li, 2023) searches the sequence space using generative models like VAE (Kingma & Welling, 2014). While these approaches allow for the creation of novel sequences, as sequences are generated from scratch, they are relatively prone to incomplete proxy models, particularly in regions where the proxy is misclassified due to being out-of-distribution.
@@
An alternative line of research is evolutionary search (Arnold, 1998; Bloom & Arnold, 2009; Schreiber et al., 2020; Sinai et al., 2020; Ren et al., 2022; Ghari et al., 2023; Kirjner et al., 2024), a popular method in biological sequence design. Especially Ghari et al. (2023) proposed GFNSeqEditor, which utilizes GFlowNets as prior distribution to edit biological sequences as an evolutionary search. Evolutionary search methods iteratively edit given sequences and constrain the new sequences so as not to deviate too far from the seed sequence; they usually start from the wild-type, which occurs in nature. This can be viewed as constrained optimization, where out-of-distribution for the proxy model can lead to unrealistic and low-score biological sequences. Consequently, they do not aim to produce highly novel sequences. Our method can be seen as a hybrid of off-policy RL and evolutionary search, capitalizing on both the high novelty offered by GFlowNets and the high rewards with out-of-distribution robustness provided by constrained search where they are properly balanced by $\delta$. Our experimental comparison with GFNSeqEditor (Ghari et al., 2023) demonstrates this hybridization balanced by $\delta$ enables us to discover sequences with greater novelty and higher rewards rather than merely using the GFlowNet policy as editing priors.
@@ 6 Experiments
Following the FLEXS benchmark (Sinai et al., 2020), [^1] FLEXS (Fitness Landscape EXploration Sandbox) is a widely-used open-source simulation environment for biological sequence design, which is available at https://github.com/samsinai/FLEXS. we evaluate our proposed method on various biological sequence design tasks. Furthermore, we analyze the effect of $\delta$-CS by directly comparing with GFN-AL on TF-Bind-8 and an anti-microbial peptide design in Section 6.6. For each experiment, we conduct five independent runs.
Implementation details. For proxy models, we employ a convolutional neural network (CNN) with one-dimensional convolutions (Sinai et al., 2020) with a UCB acquisition function and an ensemble of three network instances to measure the uncertainty. Note that we use the same architecture to implement proxy models for all baselines. For the GFlowNet policy, we use a simple two-layer long short-term memory (LSTM) network (Hochreiter & Schmidhuber, 1997) and train the policy with 1,000 inner-loop updates using a learning rate of $5\times 10^{-4}$ with a batch size of 256. However, in TF-Bind-8 and AMP, where we analyze the effectiveness of $\delta$-CS compared to GFN-AL, we directly implement $\delta$-CS on top of the GFN-AL implementation. Lastly, we set $\delta$ and $\lambda$ according to the description in Section 4; specifically, $\delta=0.5$ for tasks with $L\leq 50$ and $\delta=0.05$ for long-sequences. More details are provided in Section A.2.
@@
Baselines. As our baseline methods, we employ representative exploration algorithms. Further details are provided in Section A.3.
@@
  • AdaLead (Sinai et al., 2020) is a well-implemented model-guided evaluation method with a hill-climbing algorithm.
@@
  • CMA-ES (Hansen, 2006) is another well-known evolutionary algorithm that optimizes a continuous relaxation of one-hot vectors encoding sequence with the covariance matrix.
@@
  • GFN-AL (Jain et al., ) is our main baseline that uses GFN with Bayesian active learning.
Table 1: Maximum rewards achieved by each baseline method across six tasks, with the sequence length ($L$) for each task specified. The mean and standard deviation from five runs are reported. The highest values for each task are highlighted in bold.
RNA-A ($L=14$) RNA-B ($L=14$) RNA-C ($L=14$) TF-Bind-8 ($L=8$) GFP ($L=238$) AAV ($L=90$)
AdaLead 0.968 $\pm$ 0.070 0.965 $\pm$ 0.033 0.723 $\pm$ 0.057 0.995 $\pm$ 0.004 3.581 $\pm$ 0.004 0.565 $\pm$ 0.027
BO 0.722 $\pm$ 0.025 0.720 $\pm$ 0.032 0.506 $\pm$ 0.003 0.977 $\pm$ 0.008 3.572 $\pm$ 0.000 0.500 $\pm$ 0.000
CMA-ES 0.816 $\pm$ 0.030 0.850 $\pm$ 0.063 0.496 $\pm$ 0.041 0.986 $\pm$ 0.008 3.572 $\pm$ 0.000 0.500 $\pm$ 0.000
CbAS 0.678 $\pm$ 0.020 0.668 $\pm$ 0.021 0.492 $\pm$ 0.018 0.988 $\pm$ 0.004 3.572 $\pm$ 0.000 0.500 $\pm$ 0.000
DbAS 0.670 $\pm$ 0.041 0.652 $\pm$ 0.021 0.495 $\pm$ 0.010 0.987 $\pm$ 0.004 3.572 $\pm$ 0.000 0.500 $\pm$ 0.000
DyNA PPO 0.737 $\pm$ 0.022 0.730 $\pm$ 0.088 0.478 $\pm$ 0.015 0.977 $\pm$ 0.013 3.572 $\pm$ 0.000 0.500 $\pm$ 0.000
GFN-AL 1.030 $\pm$ 0.024 1.001 $\pm$ 0.016 0.774 $\pm$ 0.004 0.976 $\pm$ 0.002 3.578 $\pm$ 0.003 0.560 $\pm$ 0.008
GFN-AL + $\delta$-CS 1.055 $\pm$ 0.000 1.014 $\pm$ 0.001 0.972 $\pm$ 0.043 0.981 $\pm$ 0.002 3.592 $\pm$ 0.003 0.708 $\pm$ 0.010
Figure 2: Median scores of Top-128 over active rounds. Ours (GFN-AL + $\delta$-CS) consistently outperforms baseline in RNA, DNA (TF-Bind-8), and protein (GFP and AAV) design tasks.
For each task, we conduct 10 active learning rounds starting from the initial dataset $\mathcal{D}_{0}$. The query batch size is all set as 128 except for the AMP design, whose query size is 1,000. Further details of each task are provided in the following subsections. To evaluate the performance, we measure the maximum, median, and mean scores of Top-$K$ sequences.
6.1 RNA sequence design
Task setup. The goal is to design an RNA sequence that binds to the target with the lowest binding energy, which is measured by ViennaRNA (Lorenz et al., 2011). The length ($L$) of RNA is set to 14, with 4 tokens. In this paper, we have three RNA binding tasks, RNA-A, RNA-B, and RNA-C, whose initial datasets consist of 5,000 randomly generated sequences with certain thresholds; we adopt the offline dataset provided in Kim et al. (2023). We use $\delta=0.5$ and $\lambda=5$, according to the guidelines in Section 4.
Results. As shown in Figure 2 and Table 1, our method outperforms all baseline approaches. The curve in Figure 2 increases significantly faster than the other methods, indicating that $\delta$-CS effectively trains the policy and generates appropriate queries in each active round. More results are provided in Section C.1.
6.2 DNA sequence design
Task setup. In this task, we aim to generate diverse and novel DNA sequences that maximize the binding affinity to the target transcription factor. The length $(L)$ of the sequence is fixed with 8. The initial dataset $\mathcal{D}_{0}$ is the bottom 50% in terms of the reward, which results in $32,898$ samples, with the maximum score of 0.439. Though this has been widely used in many studies (Sinai et al., 2020; Jain et al., ; Trabucco et al., 2022; Kim et al., 2023), the TF-Bind-8 is easy to optimize, especially due to its size (Sinai et al., 2020). Similar to RNA, we use $\delta=0.5$ and $\lambda=5$.
Results. As shown in Table 1, AdaLead achieves the highest maximum performance, while $\delta$-CS still outperforms GFN-AL. We believe that AdaLead’s greedy evolutionary search capability is powerful, especially in the small search space of TF-bind-8. However, in Figure 2, $\delta$-CS demonstrates the best median performance compared to the other baselines; see the mean, diversity, and novelty in Section C.2.
6.3 Protein sequence design
@@
Results. Table 1 shows the results of all methods in protein sequence design tasks. Given the combinatorially vast design space with sequence lengths $L=238$ and $90$, most baselines fail to discover new sequences whose score is higher than the maximum of the dataset. In contrast, as depicted in Figure 2, our method finds high-score sequences beyond the dataset, even with a single active round. This underscores the superiority of our search strategy in practical biological sequence design tasks. Full results are provided in Section C.3.
6.4 Anti-microbial peptide design

### v2

@@ 5.1 Biological Sequence Design
Designing biological sequences using machine learning methods is widely studied. Bayesian optimization (BO) methods (Mockus, 2005; Belanger et al., 2019; Zhang et al., 2022) exploit posterior inference over newly acquired data points to update a Bayesian proxy model that can measure useful uncertainty. The BO method can be greatly improved in high-dimensional tasks by using trust-region-based search restrictions (Wan et al., 2021; Eriksson et al., 2019; Biswas et al., 2021; Khan et al., 2023) and by combining it with deep generative models (Stanton et al., 2022; Gruver et al., 2024). However, these methods usually suffer from scalability issues due to the complexity of the Gaussian process (GP) kernel (Belanger et al., 2019) or the difficulty of sampling from an intractable posterior (Zhang et al., 2022).
@@
Offline model-based optimization (MBO) (Kumar & Levine, 2020; Trabucco et al., 2021; Yu et al., 2021; Chen et al., 2022; Kim et al., 2023; Chen et al., 2023a; Yun et al., 2024) also addresses the design of biological sequences using offline datasets only, which can be highly efficient because they do not require oracle queries. These approaches have reported meaningful findings, such as the conservative requirements on proxy models since proxy models tend to give high rewards on unseen samples (Trabucco et al., 2021; Yu et al., 2021; Yuan et al., 2023; Chen et al., 2023b). This supports our approach of adaptive conservatism in the search process. Surana et al. (2024) recently noted that offline design and existing benchmarks are insufficient to reflect biological reliability, indicating that settings without additional oracle queries might be too idealistic. For a more comprehensive discussion, see the work by Kim et al. (2025).
@@
Reinforcement learning methods, such as DyNA PPO (Angermueller et al., 2020) and GFlowNets (Bengio et al., 2021; Jain et al., 2022, 2023b; Hernández-García et al., 2024), and sampling with generative models (Brookes & Listgarten, 2018; Brookes et al., 2019; Das et al., 2021; Song & Li, 2023) aim to search the biological sequence space using a sequential decision-making process with a policy, starting from scratch. Similarly, sampling with generative models (Brookes & Listgarten, 2018; Brookes et al., 2019; Song & Li, 2023) searches the sequence space using generative models like VAE (Kingma & Welling, 2014). While these approaches allow for the creation of novel sequences, as sequences are generated from scratch, they are relatively prone to incomplete proxy models, particularly in regions where the proxy is misspecified due to being out-of-distribution.
@@
An alternative line of research is evolutionary search (Arnold, 1998; Bloom & Arnold, 2009; Schreiber et al., 2020; Sinai et al., 2020; Ren et al., 2022; Ghari et al., 2023; Kirjner et al., 2024), a popular method in biological sequence design. These methods iteratively edit sequences and constrain new sequences so as not to deviate too far from the seed sequence; usually the wild-type, which occurs in nature. This can be viewed as constrained optimization, where out-of-distribution inputs to the proxy model get avoided, as they can lead to unrealistic and low-score biological sequences. As such, they do not aim to produce highly novel sequences.
5.2 GFlowNets
GFlowNets were introduced by Bengio et al. (2021) and unified by Bengio et al. (2023), demonstrating effectiveness across various domains, including language modeling (Hu et al., 2023), diffusion models (Sendera et al., 2024; Venkatraman et al., 2024), and scientific discovery (Jain et al., 2022, 2023a; AI4Science et al., 2023). Several works have aimed to improve their training methods for better credit assignment (Malkin et al., 2022; Madan et al., 2023; Pan et al., 2023; Jang et al., 2024) and extensions to multi-objective settings (Jain et al., 2023b; Chen & Mauch, 2024). Complementary to these, several studies have investigated advanced off-policy exploration strategies (Rector-Brooks et al., 2023; Shen et al., 2023; Kim et al., 2024d, c, a, b; Madan et al., 2025). Our method is particularly related to these exploration methods, yet the major difference is that they are designed under the assumption that the reward model is accurate, which does not hold in active learning and thus requires conservativeness.
GFlowNets for biological sequence design. GFlowNets were first introduced for active learning in biological sequence design by Jain et al. (2022), showing that off-policy updates can improve training stability and efficiency by leveraging static annotated datasets. They also incorporated epistemic uncertainty to encourage exploration in a curiosity-driven manner. Separately, Ghari et al. (2023) proposed GFNSeqEditor, which uses GFlowNets as priors for evolutionary editing of biological sequences, rather than for de novo generation. Their flow model is trained offline (i.e., without a proxy model) and used to generate diverse edits for a given input sequence. Our method can be seen as a hybrid of off-policy RL and evolutionary search, capitalizing on both the high novelty offered by GFlowNets and the high rewards with out-of-distribution robustness provided by constrained search, when they are properly balanced by $\delta$. Our experimental results in Section 6.3 demonstrate that $\delta$-CS achieves balanced exploration.
In this paper, we focus on applying $\delta$-CS within GFlowNets. Recent work has established a theoretical equivalence between GFlowNet training objectives, such as detailed balance and trajectory balance, and well-known off-policy maximum entropy RL algorithms, including Soft Q-Learning (Haarnoja et al., 2017) and Path Consistency Learning (Nachum et al., 2017), when appropriate reward corrections are applied (Tiapkin et al., 2024; Deleu et al., 2024). The experimental results are also provided in Section B.2.
@@ 6 Experiments
We first investigate proxy failures and the impact of our conservative search by directly applying $\delta$-CS to GFN-AL (Jain et al., 2022) in Section 6.1. Subsequently, in Section 6.2, we evaluate the proposed method on tasks from the FLEXS benchmark (Sinai et al., 2020). Section 6.3 provides a deeper comparison with GFNSeqEditor and GFN-AL, highlighting $\delta$-CS’s balanced use of proxy models. Finally, we investigate noise injection and denoising policies by comparing with local search GflowNets (Kim et al., 2024d) in Section 6.4. All results are reported over five independent runs. [^2] Available at https://github.com/hyeonahkimm/delta_cs.
6.1 Study on proxy failure and the effect of $\delta$-CS
Task: Hard TF-Bind-8. We aim to generate DNA sequences (length $L=8$) that maximize the binding affinity to the target transcription factor. Comprehensive analysis is allowed since the full sequence space is characterized by experiments (Barrera et al., 2016). However, the original task, especially due to the large initial dataset, is easy to optimize (Sinai et al., 2020), making performance differences among methods less clear. To clearly verify the effect of $\delta$-CS, we introduce a hard variant by modifying both the initial dataset distribution and the scoring landscape. Specifically, we collect the initial dataset near a certain sequence while ensuring that the initial sequences have lower scores than the given sequence to form a dataset $\mathcal{D}_{0}$ of size 1,024. We further modify the landscape to assign zero to any sequence scoring below 0.3. These modifications better reflect real-world design scenarios where the search space is vast, data is limited, and many sequences can have near-zero scores.
Dist. from $\mathcal{D}_{0}$ $\leq 0$ $\leq 1$ $\leq 2$ $\leq 8$
Spearman $\rho(f,f_{\phi})$ 0.893 0.799 0.278 0.202
Figure 2: Proxy failure on Hard TF-Bind-8. ‘$\leq 0$’ and ‘$\leq 8$’ denote the initial dataset and the whole sequence space, respectively.
Figure 3: Mean score over rounds on Hard TF-Bind-8.
Proxy failure. Figure 2 illustrates the proxy values and true scores for all $x\in\mathcal{X}$ in the first round. While the proxy provides accurate predictions for the initial data points $x\in\mathcal{D}_{0}$ (represented by the red dots), it produces unreliable predictions for points outside $\mathcal{D}_{0}$. The correlation between $f$ and $f_{\phi}$ significantly decreases as data points move farther from the observed initial dataset (denoted as ‘$\leq 0$’ in the table). This supports our hypothesis that the proxy model performs poorly on out-of-distribution data.
Figure 4: Mean scores of Top-128 over active rounds. Ours (GFN-AL + $\delta$-CS) consistently outperforms baseline in RNA, DNA (TF-Bind-8), and protein (GFP and AAV) design tasks.
Effect of $\delta$ conservativeness. As shown in Figure 3, $\delta$-CS with $\delta<1$ significantly improves GFN-AL by focusing on sequences that maintain higher correlation with the ground-truth scores. By restricting the search space to a smaller edit distance, $\delta$-CS better aligns with ground-truth scores. By limiting the search within these constrained edit distances, $\delta$-CS enhances the correlation with the oracle. Even with a fixed $\delta$, Figure 9 in Section B.6 shows consistent gains. We also confirm similar improvements on the anti-microbial peptide design (AMP) task from Jain et al. (2022); detailed results in Section B.1 show that $\delta$-CS consistently outperforms the original GFN-AL.
6.2 Experiments on FLEXS Benchmark
Implementation details. For proxy models, we employ a convolutional neural network with one-dimensional convolutions according to (Sinai et al., 2020). We use a UCB acquisition function and measure the uncertainty with an ensemble of three network instances. We use $\delta_{\text{const}}=0.5$ for DNA ($L=8$) and RNA ($L=14$) sequence design and $\delta=0.05$ for protein design ($L=238,90$). Lastly, we set $\lambda$ to satisfy $\lambda\mathbb{E}_{{\mathcal{D}_{0}}(x)}{\sigma(x)}\approx\frac{1}{L}$ based on the observations from the initial round. More details are provided in Section A.1 and Section A.2.
@@ 6.2 Experiments on FLEXS Benchmark
Baselines. As our baseline methods, we employ representative exploration algorithms. We use the same architecture to implement proxy models for all baselines except for GFN-AL, where we adopt the original implementation. Further details are provided in Section A.3.
@@
  • AdaLead (Sinai et al., 2020) is a model-guided evaluation method with a hill-climbing algorithm.
@@
  • TuRBO (Eriksson et al., 2019) is a Bayesian optimization algorithm that partitions search space into promising local regions and performs the optimization process within the region.
  • CMA-ES (Hansen, 2006) is a well-known evolutionary algorithm that optimizes a continuous relaxation of one-hot vectors encoding sequence with the covariance matrix.
@@
  • GFN-AL (Jain et al., 2022) is our main baseline that uses GFN with Bayesian active learning.
Experiment setup. For each task, we conduct 10 active learning rounds starting from the initial dataset $\mathcal{D}_{0}$. The query batch size is all set as 128. Further details of each task are provided in the following subsections. To evaluate the performance, we measure the maximum, median, and mean scores of Top-$K$ sequences.
6.2.1 RNA Sequence Design
Task. The goal is to design an RNA sequence that binds to the target with the lowest binding energy, which is measured by ViennaRNA (Lorenz et al., 2011). The length ($L$) of RNA is set to 14, with 4 tokens. In this paper, we have three RNA binding tasks, RNA-A, RNA-B, and RNA-C, whose initial datasets consist of 5,000 randomly generated sequences with certain thresholds; we adopt the offline dataset provided in Kim et al. (2023). We use $\delta=0.5$ and $\lambda=5$.
Results. As shown in Figure 4, our method outperforms all baseline approaches. The curve in Figure 4 increases significantly faster than the other methods, indicating that $\delta$-CS effectively trains the policy and generates appropriate queries in each active round. See Section C.1 in details.
6.2.2 DNA Sequence Design
Task. In this task, we follow the setup from the previous studies (Sinai et al., 2020; Jain et al., 2022; Trabucco et al., 2022; Kim et al., 2023). The initial dataset $\mathcal{D}_{0}$ is the bottom 50% in terms of the score, which results in $32,898$ samples, with a maximum score of 0.439. Similar to RNA, we use $\delta=0.5$ and $\lambda=5$.
Results. In TF-Bind-8, every method successively discovers DNA sequences with high binding affinity scores in the early rounds. Nevertheless, the results show that $\delta$-CS further improves GFN-AL by restricting search space; see the max, median, diversity, and novelty in Section C.2. In addition, $\delta$-CS demonstrates the best median scores, as well.
6.2.3 Protein Sequence Design
@@
Results. Figure 4 shows the results of all methods in protein sequence design tasks. Given the combinatorially vast design space with sequence lengths $L=238$ and $90$, most baselines fail to discover new sequences whose score is higher than the maximum of the dataset. In contrast, our method finds high-score sequences beyond the dataset, even with a single active round. This underscores the superiority of our search strategy in practical biological sequence design tasks. Full results are provided in Section C.3.
6.3 Achieving Pareto Frontier with Balanced Conservativeness
In this analysis, we demonstrate that $\delta$-CS achieves a balanced search using $\delta$, producing Pareto frontiers or comparable results to the baseline methods: GFN-AL (Jain et al., 2022) and GFNSeqEditor (Ghari et al., 2023). Notably, GFN-AL can be seen as a variant of our method with $\delta=1$, which fully utilizes the entire trajectory search. This approach is expected to yield high novelty and diversity, but it is also prone to generating low rewards due to the increased risk of out-of-distribution samples affecting the proxy model. GFNSeqEditor, on the other hand, leverages GFlowNets as a prior, editing from a wild-type sequence. It is designed to deliver reliable performance and be more robust to out-of-distribution issues by constraining the search to sequences similar to the wild type. However, unlike $\delta$-CS, GFNSeqEditor does not utilize such obtained samples for training GFlowNets in full trajectory level; GFNSeqEditor is expected to have lower diversity and novelty compared to GFN-AL and $\delta$-CS.
As shown in Figure 5, GFN-AL generally produces higher diversity and novelty in the RNA and GFP tasks compared to GFNSeqEditor. However, GFNSeqEditor performs better in terms of reward on the large-scale GFP task, whereas GFN-AL struggles due to the lack of a constrained search procedure in such a large combinatorial space. In contrast, $\delta$-CS achieves Pareto-optimal performance compared to both methods, clearly outperforming GFNSeqEditor across six tasks, with higher rewards, diversity, and novelty. For the RNA and GFP tasks, we achieve higher scores than GFN-AL while maintaining similar novelty but slightly lower diversity. In the AAV task, $\delta$-CS shows a distinct Pareto improvement. These results demonstrate that $\delta$-CS provides a beneficial balance by combining conservative search with amortized inference on full trajectories using off-policy GFlowNets training, effectively capturing the strengths of both GFN-AL and GFNSeqEditor. The results with various $\delta$ are provided in Section B.8.
6.4 Comparison with Back-and-Forth Search
(a) RNA-A
(b) GFP
Figure 6: Comparison $\delta$-CS with back-and-forth search in LS-GFN. Further results are provided in Section B.3
The noise injection and denoising policy shares similarities with the back-and-forth strategy suggested in local search GFlowNets (LS-GFN; Kim et al., 2024d). LS-GFN uses a backward policy to partially destroy a solution and a forward policy to reconstruct a new solution. While LS-GFN does not inherently use proxy models, we adapt it for a conservative search comparison by setting its backtracking step $K$ to $\delta L$, aligning its level of conservativeness with ours.
As illustrated in Figure 6, our method achieves more balanced trade-offs among high scores, diversity, and novelty than LS-GFN, likely due to its flexibility in selecting which tokens to mask. In an auto-regressive sequence generation setting, the backtracking in LS-GFN destroys only the last tokens, limiting the local search region, whereas our approach can randomly mask any token in a sequence. Moreover, as shown in Table 3 and Figure 8, our approach demonstrates it remains more robust when the initial dataset is restricted to lower-quality sequences; see Section B.4.
6.5 Further Studies
Studies on the effect of $\delta$-CS with various off-policy RL.
Studies on $\delta$. We study the choice of $\delta_{\text{const}}$ and the effectiveness of the adaptive $\delta(x,\sigma)$. In Section B.7 and Section B.8, we investigate how varying $\delta_{\text{const}}$ affects performance. While $\delta$-CS generally outperforms GFN-AL, its gains diminish as $\delta_{\text{const}}$ grows, because $\delta=1$ reverts to on-policy exploration without any conservative constraint. Furthermore, the results in Section B.6 verify the benefit of adaptive $\delta(x,\sigma)$. While even a fixed $\delta$ improves GFN-AL on Hard TF-Bind-8, adaptive $\delta$ yields additional performance gains in overall.
Studies on the varying proxy qualities. We further validate $\delta$-CS under degraded proxy models by truncating the initial dataset to its 50%, 25%, and 10% score percentiles. As shown in Figure 8, across all truncation levels, $\delta$-CS consistently outperforms GFN-AL and LS-GFN, confirming that incorporating conservative search is crucial for robustness against proxy misspecification; the details are provided in Section B.4.
7 Conclusions and discussion
Limitations and future works. The main limitation of our method is that it does not fundamentally resolve the drawbacks of active learning; it serves as a useful tool within the existing framework. Investigating robust proxy training and uncertainty measurement techniques remains necessary. These improvements are orthogonal to our approach and can enhance $\delta$-CS when integrated. Future work includes combining $\delta$-CS with existing GFlowNet methods. For example, applying improved credit assignment for larger-scale tasks (Jang et al., 2024) and extending to multi-objective settings (Jain et al., 2023b; Chen & Mauch, 2024) could significantly boost its applicability and effectiveness.
Impact Statement
This work tackles a common challenge in scientific discovery: generating candidates with limited data and unreliable proxy models. In biological sequence design, we introduce a conservative search strategy that adapts exploration range based on proxy uncertainty to reduce reward hacking. While developed for this domain, the principle of aligning exploration with model confidence may generalize to other domains where generative models and active learning pipelines face similar constraints.
Acknowledgements
Minsu Kim is supported by KAIST Jang Yeong Sil Fellowship and the Canadian AI Safety Institute Research Program at CIFAR through a Catalyst award. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2025-00563763).
Appendix A Implementation detail
Algorithm 1 Active Learning GFlowNets with $\delta$-CS
2:  for $t=1$ to $T$ do
[…]
4:     for $k=1$ to numProxyTrain do
5:        train proxy $f_{\phi}(x)$ with current round dataset $\mathcal{D}_{t-1}$: $\mathcal{L}(\phi)=\mathbb{E}_{x\sim P_{\mathcal{D}_{t-1}}(x)}\left[\left(f(x)-f_{\phi}(x)\right)^{2}\right].$
6:     end for
[…]
8:     for $k=1$ to numPolicyTrain do
9:        obtain off-policy trajectories $\hat{\tau}_{1},\ldots,\hat{\tau}_{M}$ from $\hat{x}_{1},\ldots,\hat{x}_{M}$ given by $\delta$-CS ($\mathcal{D}_{t-1},M,\delta$).
10:        obtain offline trajectory $\tau_{1},\ldots,\tau_{M}$ from $x_{1},\ldots,x_{M}\sim P_{\mathcal{D}_{t-1}}(\tau)$.
11:        train $\theta$ with TB loss over $\hat{\tau}_{1},\ldots,\hat{\tau}_{M}$ and $\tau_{1},\ldots,\tau_{M}$
$\mathcal{L}(\theta)=\frac{1}{2M}\sum_{i=1}^{M}\left(\log\frac{Z_{\theta}P_{F}(\tau_{i};\theta)}{R(x_{i};\phi)}\right)^{2}+\frac{1}{2M}\sum_{i=1}^{M}\left(\log\frac{Z_{\theta}P_{F}(\hat{\tau}_{i};\theta)}{R(\hat{x}_{i};\phi)}\right)^{2}.$
12:     end for
[…]
14:     obtain query samples $\hat{x}_{1},\ldots,\hat{x}_{B}$ from $\delta$-CS ($\mathcal{D}_{t-1}$, $B$, $\delta$).
15:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{(\hat{x}_{i},f(\hat{x}_{i}))\}_{i=1}^{B}$.
16:  end for
Algorithm 2 Sampling with $\delta$-CS
1:  Input: Dataset $\mathcal{D}_{t-1}$, batch size $M$, conservative parameter $\delta$
2:  sample high reward data $x_{1},\ldots,x_{M}$ with rank-based reweighed prior $P_{\mathcal{D}_{t-1}}(\cdot)$.
3:  obtain masked data $\tilde{x}_{1},\ldots,\tilde{x}_{M}$ with noise injection policy $P_{\text{noise}}(\tilde{\cdot}|\cdot,\delta)$ from $x_{1},\ldots,x_{M}$.
4:  obtain denoised $\hat{x}_{1},\ldots,\hat{x}_{M}$ with denoising policy $P_{\text{denoise}}(\hat{\cdot}\mid\tilde{\cdot};\theta)$ from $\tilde{x}_{1},\ldots,\tilde{x}_{M}$.
5:  return $\hat{x}_{1},\ldots,\hat{x}_{M}$.
A.1 Proxy training
For training proxy models, we follow the procedure of (Jain et al., 2022). We use Adam (Kingma, 2015) optimizer with learning rate $1\times 10^{-5}$ and batch size of 256. The maximum proxy update is set as 3000. To prevent over-fitting, we use early stopping using the $10\%$ of the dataset as a validation set and terminate the training procedure if validation loss does not improve for five consecutive iterations.
A.2 Policy training
As described in Section 6, we employ a two-layer long short-term memory (LSTM; Hochreiter & Schmidhuber, 1997) with 512 hidden dimensions. The policy is trained with a learning rate of $5\times 10^{-4}$ with a batch size of 256. The learning rate of $Z$ is set as $10^{-3}$. The coefficient $\kappa$ in Section 3.2 is set as 0.1 for TF-Bind-8 and AMP with MC dropout, according to Jain et al. (2022), and 1.0 for RNA and protein design with Ensemble following Ren et al. (2022).
A.3 Implementation details of baselines
We adopt open-source code from FLEXS benchmark (Sinai et al., 2020).
  • TuRBO (Eriksson et al., 2019): We closely follow the algorithm presented in (Eriksson et al., 2019). As it supports continuous search space, convert the continuous vectors into one-hot representation by computing the argmax at each sequence position.
  • GFN-AL (Jain et al., 2022): We strictly follow hyperparameters of the original code in they conduct experiments on TF-Bind-8 and AMP. The proxy is parameterized using an MLP with two layers of 2,048 hidden. For the policy, a 2-layer MLP with 2,048 hidden dimensions is used, but we also test it with a 2-layer LSTM policy.
Appendix B Further studies
B.1 Anti-microbial peptide design
Task setup. The goal is to generate protein sequences with anti-microbial properties (AMP). The vocabulary size $|\mathcal{V}|=20$, and the sequence length ($L$) varies across sequences, and we consider sequences of length 50 or lower. For the AMP task, we consider a much larger query batch size for each active round because they can be efficiently synthesized and evaluated (Jain et al., 2022). We set $\delta$ as 0.5 with $\lambda=1$.